# Subsystem localization

### v2

@@ III Static properties
To analyze the localization character of the $\epsilon$-significant eigenkets, for example, to see whether an $\epsilon$-significant eigenket is localized within the subsystem $A$ or not, we now compute the inverse participation ratio $IPR_{A}$ (as in Eq. 3) for these states. Fig. 4 shows the variation of $IPR_{A}$ as a function of $V$ for different values of $\epsilon$, specifically $\epsilon$ = 0.3, 0.5, and 0.7. These calculations are performed for a total system size 2N = 1000, considering three different values of $t^{\prime}$, namely $t^{\prime}$ = 1, 5, and 10. The results clearly indicate that for small $t^{\prime}$ (in particular, $t^{\prime}=1$), most of the $\epsilon$-significant eigenkets are not localized within the subsystem $A$. As we increase $t^{\prime}$ beyond a critical value $t^{\prime}_{c}$, we see the appearance of a localized phase for intermediate values of $V$. For a give $t^{\prime}>t^{\prime}_{c}$, and for $V_{1}<V<V_{2}$, almost all the $\epsilon$-significant eigenkets are localized up to a good extent but the states in the middle of the spectrum show stronger localization. For a given $t^{\prime}$, we calculate $V_{1}$ and $V_{2}$ in the following way. For a state in the middle of the spectrum, we choose $V_{1}$ in such a way that $IPR_{A}<0.5$ for $V<V_{1}$ and $IPR_{A}\geq 0.5$ for $V\geq V_{1}$. Similarly, for the same state, we choose $V_{2}$ in such a way that $IPR_{A}>0.5$ for $V<V_{2}$ and $IPR_{A}\leq 0.5$ for $V\geq V_{2}$. The plots for $V_{1}$ and $V_{2}$ as functions of $t^{\prime}$ can be seen in Fig. 5. These numerical results also help us find $t^{\prime}_{c}$ by extrapolating $V_{1}$ and $V_{2}$ till they become equal. We estimate that $t^{\prime}_{c}\approx 4.6$. It is interesting to note that, as $t^{\prime}$ increases beyond the critical value $t^{\prime}_{c}$, the value of $V_{2}$ increases, while the value of $V_{1}$ decreases. The decrease in $V_{1}$ with increasing $t^{\prime}$ can be understood from the fact that, in the limit of very large $t^{\prime}$, the subsystem $A$ tends toward a fully localized phase, since the eigenstates are localized along the rungs of the ladder, for any finite $V$.
@@
The analysis of the subsystem for $V>V_{2}$ (and $t^{\prime}>t^{\prime}_{c}$) reveals a new phase which we call a weakly delocalized phase with the fractal character. In this parameter regime, some states exhibit delocalization properties (having small $IPR_{A}$), while other states (near the edges) are found to be well localized (having large $IPR_{A}$). Such reentrance of the delocalized phase at large values of the quasiperiodic potential strength has recently been observed in several one-dimensional models. [47, 48]. To better understand the localization properties of different phases of the subsystem, we next perform the finite size scaling of the average participation ratio, $\langle PR_{A}\rangle$, where the average is taken only over the $\epsilon$-significant eigenkets:
Here $\mathcal{S}_{\epsilon}$ is the set of the $\epsilon$-significant eigenkets (also see Eq. 5). Fig. 6 shows how $\langle PR_{A}\rangle$ scales with the subsystem size (N) for $V$ = 1, 10 and 17 (for $t^{\prime}$ = 5 and $\epsilon$ = 0.5). Corresponding log-log plots, as seen in Fig. 7, demonstrate that the quantity of interest obeys $\langle PR_{A}\rangle\sim N^{\eta}$ scaling in general. We find that $\eta=1$ when $V=1$ (i.e., when $V<V_{1}$). This result confirms that the subsystem is in delocalized phase in the regime where $V$ is small. For $V=10$ (i.e., when $V_{1}<V<V_{2}$), we find $\eta\approx 0$. This shows that the subsystem $A$ is in a localized phase in the intermediate range of $V$. When $V=17$ (i.e., when $V>V_{2}$), we get $\eta<1$. This indicates that the corresponding phase is a fractal. Interestingly, when $V=5$ (i.e., when $V$ is close to $V_{1}$), we also get $\eta<1$. This observation suggests the presence of a narrow band along the transition line between the delocalized and localized phases, within which the subsystem exhibits a fractal phase.
When $t^{\prime}<t^{\prime}_{c}$, the similar analysis of $IPR_{A}$ and $\langle PR_{A}\rangle$ shows that there is no localized phase and we have only two phases - a delocalized phase and weakly delocalized (or fractal) phase. The relevant results may be found in Figs. 15 and 16.
For better understanding of these fractal phases (and other two phases, namely, delocalized and localized phases), we next study the wavepacket dynamics in the subsystem.
@@ IV Dynamic properties
Figure 9: $\sigma^{2}$ vs $t$ plot for different $V$ for subsystem size $N=1000$. The black dashed lines show the fitting to the functional form $\sigma^{2}\sim t^{\gamma}$. The inset shows the variation of the exponent $\gamma$ on $V$.
Figure 10: Variation of $P(l)$, the Probability of finding the particle in the region $N/2-l$ to $N/2+l$, with time $t$ for different $V$. Results are for $N=1000$ and $l=10$. Plots are generated using local time averaging to smooth out the data.
@@
Here, we first focus on $t^{\prime}=5>t_{c}^{\prime}$. Figure 8 shows the time evolution of the probability distribution $|\psi(j,t)|^{2}=|\langle j|\psi(t)\rangle|^{2}$ for $V=1$, $10$, and $17$. According to our static results, these three values of $V$ correspond to completely delocalized, completely localized, and weakly delocalized (fractal) phases, respectively. We observe that for $V=1$, $|\psi(j,t)|^{2}$ spreads significantly faster across subsystem $A$ compared to the case of $V=17$. In contrast, for $V=10$, the probability distribution remains nearly unchanged over time, indicating a fully localized phase.
Next, we quantify the spread of $|\psi(j,t)|^{2}$ using various measures. We compute the variance $\sigma^{2}$ of this probability distribution, defined as
@@
In Fig. 9, we plot $\sigma^{2}$ as a function of time $t$ for different values of $V$ (with $t^{\prime}=5$ fixed). In the completely localized phase (e.g., $V=10$), we observe that $\sigma^{2}$ remains constant over time, as expected. In contrast, in other regimes, $\sigma^{2}$ grows as $\sigma^{2}\sim t^{\gamma}$ until it saturates due to finite-size effects. For small values of $V$ (e.g., $V=1$), we find the exponent $\gamma=2$, indicating ballistic transport. As $V$ increases, we enter a super-diffusive regime where $21<\gamma<$ 2 (see results for $V=5$), followed by a completely localized phase. Upon further increasing $V$ (e.g., $V=17$), we enter a weakly delocalized (fractal) regime, as identified in our static calculations, where the dynamics becomes sub-diffusive, i.e., $\gamma<1$. The variation of the exponent $\gamma$ with $V$ is shown in the inset of Fig. 9. We have verified the robustness of the exponent $\gamma$ by varying the system size. An alternative validation of subdiffusive transport, based on the scaling of the projected probability distribution, is presented in Appendix B.
@@
Figure 11: Scaling of $P(l)_{sat}$, the long time average of $P(l)$, with inverse of the subsystem size $1/N$. The black dashed line represents the fitting of the data with the functional form $P(l)_{sat}\sim(1/N)^{\alpha}+c$. The inset shows the variation of $\alpha$ and $c$ with $V$. Results are for $202l=$ 20.
@@
where $l\ll N$. This quantity represents the total probability of finding the particle at time $t$ within a neighborhood of width $2l$ centered around its initial position. For our choice of initial state, we have $P(l,0)=1$. This measure provides insight into how much memory of the initial state is retained during the dynamics. In the completely localized phase, one expects $P(l,t)\approx 1$ for all times, provided $l$ is larger than the typical localization length. On the other hand, in a completely delocalized phase, one would expect that at sufficiently large times $|\psi(j,t)|^{2}\sim\frac{1}{N}$ for all $j$, implying that $P(l,t)$ saturates to $P(l)\approx\frac{2l}{N}$. This scales to zero in the thermodynamic limit ($N\to\infty$), indicating a complete loss of memory of the initial state. Figure 10 confirms this behavior: for $V=1$ (completely delocalized) and $V=10$ (completely localized), $P(l=10,t)$ behaves as expected—decreasing rapidly with time and saturating to a small value for $V=1$, while remaining close to $1$ for all times in the case of $V=10$. Interestingly, for $V=5$ and $V=17$, $P(l,t)$ decreases at early times but saturates to a significantly higher value than in the delocalized case ($V=1$), indicating partial memory retention.
@@
[…]
Figure 11 shows the scaling of $P(l)_{\text{sat}}$ with subsystem size $N$ for different values of $V$. We fit the data using the function $\text{const}/N^{\alpha}+c$, and the variation of the fitting parameters $\alpha$ and $c$ with $V$ is shown in the inset of Fig. 11. In the completely delocalized phase ($V=1$), we find $\alpha\approx 1$ and $c\approx 0$, as expected. In the completely localized phase, $P(l)_{\text{sat}}$ remains nearly constant with $N$. Interestingly, for $V=5$ (super-diffusive regime) and $V=17$ (sub-diffusive regime), the scaling behavior is similar, but the fitted constant $c\neq 0$ in both cases. This indicates that a finite portion of the initial state’s memory survives even in the thermodynamic limit. We investigate $t^{\prime}=3<t^{\prime}_{c}$ case in the Appendix. C, and find no signature of the localized phase; instead the dynamics suggest a ballistic (for small $V$) and a subdiffusive phase (for large $V$), with an intermediate crossover regime which can be either super- or subdiffusive in nature.

### v3

@@ III.2 $IPR_{A}$: Inverse Participation Ratio for subsystem
To analyze the localization character of the $\epsilon$-significant eigenkets, we now compute $IPR_{A}$ for these states. Fig. 4 shows the dependence of $IPR_{A}$ on $V$ for the $\epsilon$-significant eigenkets with $\epsilon$ = 0.3, 0.5, and 0.7. These calculations are performed for a total system size 2N = 1000, considering three different values of $t^{\prime}$, namely $t^{\prime}$ = 1, 5, and 10. The results clearly indicate that for small $t^{\prime}$ (in particular, $t^{\prime}=1$), most of the $\epsilon$-significant eigenkets are not localized within the subsystem $A$. As we increase $t^{\prime}$ beyond a critical value $t^{\prime}_{c}$, we see the appearance of a localized phase for intermediate values of $V$. In this regime, almost all the $\epsilon$-significant eigenkets are localized up to a good extent but the states in the middle of the spectrum show stronger localization.
For a given $t^{\prime}>t^{\prime}_{c}$, the subsystem shows localization when $V$ is above a lower bound (denoted by $V_{1}$) and below an upper bound (denoted by $V_{2}$). The lower and upper bounds for localization, i.e., $V_{1}$ and $V_{2}$, depend on the fixed value of $t^{\prime}$. We calculate these two bounds in the following way. For a state in the middle of the spectrum, we choose $V_{1}$ in such a way that $IPR_{A}<0.5$ for $V<V_{1}$ and $IPR_{A}\geq 0.5$ for $V\geq V_{1}$. Similarly, for the same state, we choose $V_{2}$ in such a way that $IPR_{A}>0.5$ for $V<V_{2}$ and $IPR_{A}\leq 0.5$ for $V\geq V_{2}$. The plots for $V_{1}$ and $V_{2}$ as functions of $t^{\prime}$ can be seen in Fig. 5. These numerical results also help us find $t^{\prime}_{c}$ by extrapolating $V_{1}$ and $V_{2}$ till they become equal. This extrapolation is performed by fitting the numerical data with a quadratic polynomial. From this procedure, we estimate $t^{\prime}_{c}\approx 4.4$. It is interesting to note that, as $t^{\prime}$ increases beyond the critical value $t^{\prime}_{c}$, the value of $V_{2}$ increases, while the value of $V_{1}$ decreases. The decrease in $V_{1}$ with increasing $t^{\prime}$ can be understood from the fact that, in the limit of very large $t^{\prime}$, the subsystem $A$ tends toward a fully localized phase, since the eigenstates are localized along the rungs of the ladder, for any finite $V$.
The analysis of the subsystem for $V>V_{2}$ (and $t^{\prime}>t^{\prime}_{c}$) reveals a new phase which we call a weakly delocalized phase with the fractal character. In this parameter regime, some states exhibit delocalization properties (having small $IPR_{A}$), while other states (near the edges) are found to be well localized (having large $IPR_{A}$). Such reentrance of the delocalized phase at large values of the quasiperiodic potential strength $V$ has recently been observed in several one-dimensional models [49, 50].
For completeness, we also look into the energy-resolved IPR for the subsystem $A$, bath $B$ and the full system. The results may be found in Appendix C. In analyzing the localization properties of subsystem $A$, the role of the admixture of localized and delocalized (weakly delocalized) states is explained in the same appendix.
III.3 $\langle PR_{A}\rangle$: Average Participation Ratio for subsystem
To better understand the localization properties of different phases of the subsystem, we next perform the finite size scaling of the average participation ratio, $\langle PR_{A}\rangle$, where the average is taken only over the $\epsilon$-significant eigenkets:
Here, $IPR^{(n)}_{A}$ is defined in Eq. 5 and the sum is over all the indices corresponding to the $\epsilon$-significant eigenkets in $\mathcal{S}_{\epsilon}$. We recall that $\mathcal{S}_{\epsilon}$ is the set of the $\epsilon$-significant eigenkets and $|\mathcal{S}_{\epsilon}|$ denotes the number of elements in the set (also see Eq. 3). For large subsystem size $N$, we expect the following finite size scaling law to hold:
$\langle PR_{A}\rangle\sim N^{\eta}.$ (7)
This is a special case of a more general scaling relation which can be used for studying the multifractal nature of the subsystem:
$\frac{1}{|\mathcal{S}_{\epsilon}|}\sum_{n\in\mathcal{S}_{\epsilon}}\left[\frac{1}{\sum_{j\in A}\lvert\braket{j|\psi_{n}}\rvert^{2q}}\right]\sim N^{D_{q}(q-1)}.$ (8)
Here, $D_{q}$ is the fractal dimension of the subsystem in a given phase. Our special case, as presented in Eq. 7, corresponds to $q=2$ and $D_{q}=\eta$. The subsystem is considered to be in the delocalized (localized) phase when $\eta=1$ (0). The subsystem is said to be in a fractal phase if $0<\eta<1$. We point out here that the introduction of the fractal dimension $\eta$ (or $D_{q}$) for a subsystem closely parallels its definition for the full system, where it appears in the scaling relation of the average inverse participation ratio (IPR) of the entire system [47, 50, 51].
Fig. 6 shows how $\langle PR_{A}\rangle$ scales with the subsystem size (N) for $V$ = 1, 10 and 17 (with fixed $t^{\prime}$ = 5 and $\epsilon$ = 0.5). Corresponding log-log plots, as seen in Fig. 7, demonstrate that the quantity of interest obeys $\langle PR_{A}\rangle\sim N^{\eta}$ scaling in general. We emphasize that even when the subsystem hosts both delocalized (weakly delocalized) and localized states for certain values of the model parameters, this scaling relation remains valid. Further details are provided in Appendix C.
We find that $\eta=1$ when $V=1$ (i.e., when $V<V_{1}$). This result confirms that the subsystem is in delocalized phase in the regime where $V$ is small. For $V=10$ (i.e., when $V_{1}<V<V_{2}$), we find $\eta\approx 0$. This shows that the subsystem $A$ is in a localized phase in the intermediate range of $V$. When $V=17$ (i.e., when $V>V_{2}$), we get $0<\eta<1$. This indicates that the corresponding phase is a fractal. Interestingly, when $V=5$ (i.e., when $V$ is close to $V_{1}$), we also get $0<\eta<1$. This observation suggests the presence of a narrow band along the transition line between the delocalized and localized phases, within which the subsystem exhibits a fractal phase.
III.4 Subsystem properties for $t^{\prime}<t^{\prime}_{c}$
So far we have mainly focused on the subsystem static localization properties for $t^{\prime}>t^{\prime}_{c}$. When $t^{\prime}<t^{\prime}_{c}$, the similar analysis of $IPR_{A}$ and $\langle PR_{A}\rangle$ shows that there is no localized phase and we have only two phases - a delocalized phase and a weakly delocalized (or fractal) phase. The relevant results may be found in Appendix B (especially, see Figs. 17 and 18).
Figure 8: Contour plot of the scaling exponent $\eta$ as a function of $t^{\prime}$ and V. Expected phase boundaries are indicated by white lines (see also Fig. 2).
III.5 A contour plot for scaling exponent $\eta$
To gain a clearer understanding of the static phases of subsystem $A$, we evaluate the scaling exponent $\eta$ from the expected relation $\langle PR_{A}\rangle\sim N^{\eta}$, over a carefully chosen grid in the $t^{\prime}-V$ plane. The grid is defined as follows: for $t^{\prime}$, we consider values ranging from $0.5$ to $10.0$ in steps of $0.5$ (20 values in total); for $V$, we take the set of 21 values consisting of $2.0$ to $15.0$ in steps of $1.0$, $17.0$, $19.0$, and $20.0$ to $60.0$ in steps of $10.0$. To extract the value of $\eta$, we calculate $\langle PR_{A}\rangle$ for $N=$ 300, 400, 600, 800 and 1000. The values of $\eta$ at different grid points are presented as a contour plot in Fig. 8.
This contour plot clearly shows what we expected earlier from our detailed calculations on some carefully chosen points in the parameter plane. For $t^{\prime}>t^{\prime}_{c}$, we have a localized phase ($\eta\sim 0$) in the intermediate values of the AA potential strength $V_{1}<V<V_{2}$. For $V<V_{1}$ and $V>V_{2}$, we respectively get a delocalized phase ($\eta\sim 1$) and a weakly delocalized fractal phase ($0<\eta<1$). Interestingly, we also get a narrow fractal phase along the line separating delocalized and localized phases. When $t^{\prime}<t^{\prime}_{c}$, no evidence of a localized phase is observed. For small $V$, the system remains in a delocalized phase, while for intermediate to large $V$, the subsystem enters a weakly delocalized fractal phase. Within this fractal regime, for a fixed $V$, the scaling exponent $\eta$ increases as $t^{\prime}$ decreases, and $\eta\to 1$ as $t^{\prime}\to 0$. We recall that the subsystem $A$ stays in a delocalized phase ($\eta\sim 1$) on the $t^{\prime}=0$ axis.
III.6 Understanding reentrance phenomenon
For $t^{\prime}>t^{\prime}_{c}$, the subsystem $A$ exhibits three major distinct phases as $V$ increases: a delocalized phase, followed by a localized phase, and finally a weakly delocalized (fractal) phase. This sequence represents a form of reentrance behavior, which can be understood intuitively as follows.
For small $V<2$ with $t^{\prime}=0$, both subsystem $A$ and bath $B$ are delocalized (eigenkets spread within $A$ or $B$). Introducing a finite coupling $t^{\prime}>0$ keeps both the subsystem and bath delocalized, though now the eigenkets extend across the full system $A+B$ (see Appendix C).
For $V>2$ with $t^{\prime}=0$, bath $B$ becomes localized while subsystem $A$ retains $N$ delocalized states. Increasing $t^{\prime}$ beyond a certain value then hybridizes these states in $A$ into rung-localized dimers, whose projected parts are localized in $A$ as reflected by large $IPR_{A}$.
At larger $V$ (keeping $t^{\prime}$ fixed), bath $B$ develops strongly localized states, while the remaining eigenstates in $A$ weakly couple to $B$. The projected such states form the weakly delocalized (fractal) phase.
Figure 9: Plot of $\overline{\kappa}$ as a function of V with $t^{\prime}=5$. For the calculations, we consider the subsystem size $N=500$ and take only $\epsilon=0.5$ significant states.
@@ III.6 Understanding reentrance phenomenon
[…]
To gain insight into how this bond localization parameter changes as a function of $V$ for a given $t^{\prime}$, we plot average of this parameter, defined as follows:
[…]
where $\mathcal{S}_{\epsilon}$ is the set of $\epsilon$-significant states as defined earlier (see Sec. III.1). The plot of $\overline{\kappa}$ as a function of $V$ (with $t^{\prime}=5>t^{\prime}_{c}$) can be found in Fig. 9. This result clearly demonstrates that there is indeed localization along the bonds (rungs) in the intermediate region of $V$ where we observe the subsystem localization in the reentrance phenomenon.
@@ IV Dynamic properties
Figure 11: $\sigma^{2}$ vs $t$ plot for different $V$ for subsystem size $N=1000$ and $t^{\prime}=5>t^{\prime}_{c}$. The black dashed lines show the fitting to the functional form $\sigma^{2}\sim t^{\gamma}$. The inset shows the variation of the exponent $\gamma$ on $V$.
Figure 12: Contour plot of the scaling exponent $\gamma$ as a function of $t^{\prime}$ and $V$. Expected phase boundaries are indicated by white lines (see also Fig. 2).
Figure 13: Variation of $P(l)$, the Probability of finding the particle in the region N/2lN/2-l to $N/2+l$, with time $t$ for different $V$. Results are for $N=1000$, $l=10$, and $t^{\prime}=5$. Plots are generated using local time averaging to smooth out the data.
@@
Here, we first focus on $t^{\prime}=5>t_{c}^{\prime}$. Figure 10 shows the time evolution of the probability distribution $|\psi(j,t)|^{2}=|\langle j|\psi(t)\rangle|^{2}$ for $V=1$, $10$, and $17$. According to our static results, these three values of $V$ correspond to completely delocalized, completely localized, and weakly delocalized (fractal) phases, respectively. We observe that for $V=1$, $|\psi(j,t)|^{2}$ spreads significantly faster across subsystem $A$ compared to the case of $V=17$. In contrast, for $V=10$, the probability distribution remains nearly unchanged over time, indicating a fully localized phase. Next, we quantify the spread of the probability distribution using different measures.
IV.1 MSD of the projected time-evolved state
First, we quantify the spread of the projected time-evolved state in subsystem $|\psi(j,t)|^{2}$ using the variance $\sigma^{2}$ of the probability distribution. The variance $\sigma^{2}$ of this probability distribution is defined as,
@@ IV.1 MSD of the projected time-evolved state
Typically, the long-time growth of $\sigma^{2}$ scales as
$\sigma^{2}(t)\sim t^{\gamma},$ (12)
with $0\leq\gamma\leq 2$. Ballistic transport corresponds to $\gamma=2$, while $1<\gamma<2$ ($0<\gamma<1$) indicates superdiffusive (subdiffusive) behavior. The case $\gamma=1$ represents diffusive transport, and in the localized phase one finds $\gamma=0$.
@@
In Fig. 11, we plot $\sigma^{2}$ as a function of time $t$ for different values of $V$ (with $t^{\prime}=5$ fixed). In the completely localized phase (e.g., $V=10$), we observe that $\sigma^{2}$ remains constant over time, as expected. In contrast, in other regimes, $\sigma^{2}$ grows as $\sigma^{2}\sim t^{\gamma}$ until it saturates due to finite-size effects. For small values of $V$ (e.g., $V=1$), we find the exponent $\gamma=2$, indicating ballistic transport. As $V$ increases, we enter into a super-diffusive regime where $1<\gamma<2$ (see results for $V=5$), followed by a sub-diffusive regime with $\gamma<1$ (see results for $V=6.4$), and a completely localized phase. Upon further increasing $V$ (e.g., $V=17$), we enter a weakly delocalized (fractal) regime, as identified in our static calculations, where the dynamics becomes sub-diffusive, i.e., $\gamma<1$. The variation of the exponent $\gamma$ with $V$ is shown in the inset of Fig. 11. We discuss the robustness of the exponent $\gamma$ by varying the system size in Appendix G. An alternative validation of subdiffusive transport, based on the scaling of the projected probability distribution, is presented in Appendix E.
IV.2 A contour plot for scaling exponent $\gamma$
To gain a clearer understanding of the dynamic phases of subsystem $A$, we evaluate the scaling exponent $\gamma$ from the expected relation $\sigma^{2}\sim t^{\gamma}$, over a carefully chosen grid (same as mentioned in Sec . III.5) in the $t^{\prime}-V$ plane. To obtain the exponent $\gamma$, we have fitted the long-time regime (before saturation) data of $\sigma^{2}$ vs $t$ plot for $N=1000$. The value of $\gamma$ at different grid points is presented as a contour plot in Fig. 12.
As indicated by our static calculations, for a fixed $t^{\prime}>t^{\prime}_{c}\approx 4.4$, we find that the exponent $\gamma\approx 0$ within a finite interval between $V_{1}$ and $V_{2}$, signaling the presence of a localized phase, which is identified as $P2$ phase in Fig. 2. We further note that $V_{1}$ decreases slightly while $V_{2}$ increases with increasing $t^{\prime}$. This trend suggests that in the limit $t^{\prime}\to\infty$, the subsystem $A$ becomes completely localized for any finite value of $V$, which is consistent with the intuitive picture of eigenstate dimerization in the total Hamiltonian $H$. On the other hand, the exponent $\gamma\approx 2$ for $V\ll V_{1}$, indicating a ballistic phase (which is identified as a part of $P1$ phase in Fig. 2), and $\gamma<1$ for $V>V_{2}$, suggesting a sub-diffusive phase (which is identified as a part of $P3$ phase in Fig. 2).
If $t^{\prime}<t_{c}^{\prime}$ but sufficiently large, we find a ballistic and the sub-diffusive phase in the small and the large $V$ regime, respectively (see Appendix. B for details), which is once again identified as a part of $P1$ phase and $P3$ phase in Fig. 2. If $t^{\prime}/V\ll 1$, we observe that even at large $V$ the system can exhibit a super-diffusive phase with $\gamma>1$ (which is identified as $P5$ phase in Fig. 2). Moreover, for extremely small $t^{\prime}/V$, i.e., $t^{\prime}/V\to 0$, we find $\gamma\to 2$, indicating even the existence of a ballistic phase. This ballistic phase is not visible on the scale of Fig. 12; therefore, in Appendix F, we present a zoomed-in version of the small-$t^{\prime}$ regime of the phase diagram to clarify this feature. Note that in the strict $t^{\prime}=0$ limit, the dynamics are trivially ballistic, as it is entirely governed by the Hamiltonian $H_{A}$ (see Eq. 1).
We also identify an intermediate crossover regime (which is identified as $P4$ phase in Fig. 2) between the ballistic and localized phase (for $t^{\prime}>t^{\prime}_{c}$), as well as between the ballistic and sub-diffusive phase (for $t^{\prime}<t^{\prime}_{c}$). In this phase, the exponent $\gamma$ exhibits a strong dependence on both the system size and the model parameters. Our $N=1000$ data suggests the existence of both super-diffusive and sub-diffusive phases within this crossover regime. However, a clear phase boundary between these phases is extremely difficult to establish using finite-size numerics. Further details are provided in Appendix G.
@@ IV.2 A contour plot for scaling exponent $\gamma$
Figure 14: Scaling of $P(l)_{sat}$, the long time average of $P(l)$, with inverse of the subsystem size $1/N$. The black dashed line represents the fitting of the data with the functional form $P(l)_{sat}\sim(1/N)^{\alpha}+c$. The inset shows the variation of $\alpha$ and $c$ with $V$. Results are for $2l=20$, and $t^{\prime}=5$.
IV.3 Memory of initial state
@@
where $l\ll N$. This quantity represents the total probability of finding the particle at time $t$ within a neighborhood of width $2l$ centered around its initial position. For our choice of initial state, we have $P(l,0)=1$. This measure provides insight into how much memory of the initial state is retained during the dynamics. In the completely localized phase, one expects $P(l,t)\approx 1$ for all times, provided $l$ is larger than the typical localization length. On the other hand, in a completely delocalized phase, one would expect that at sufficiently large times $|\psi(j,t)|^{2}\sim\frac{1}{N}$ for all $j$, implying that $P(l,t)$ saturates to $P(l)\approx\frac{2l}{N}$. This scales to zero in the thermodynamic limit ($N\to\infty$), indicating a complete loss of memory of the initial state. Figure 13 confirms this behavior for $t^{\prime}=5>t^{\prime}_{c}$: for $V=1$ (completely delocalized) and $V=10$ (completely localized), $P(l=10,t)$ behaves as expected—decreasing rapidly with time and saturating to a small value for $V=1$, while remaining close to $1$ for all times in the case of $V=10$. Interestingly, for $V=5$ and $V=17$, $P(l,t)$ decreases at early times but saturates to a significantly higher value than in the delocalized case ($V=1$), indicating partial memory retention.
@@
$P(l)_{\text{sat}}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}P(l,t)\,dt.$ (14)
Figure 14 shows the scaling of $P(l)_{\text{sat}}$ with subsystem size $N$ for different values of $V$ and fixed $t^{\prime}=5$. We fit the data using the function $\text{constant}/N^{\alpha}+c$, and the variation of the fitting parameters $\alpha$ and $c$ with $V$ is shown in the inset of Fig. 14. In the completely delocalized phase ($V=1$), we find $\alpha\approx 1$ and $c\approx 0$, as expected. In the completely localized phase, $P(l)_{\text{sat}}$ remains nearly constant with $N$. Interestingly, for $V=5$ (super-diffusive) and $V=17$ (sub-diffusive), the scaling behavior is similar, but the fitted constant $c\neq 0$ in both cases. This indicates that a finite portion of the initial state’s memory survives even in the thermodynamic limit.